OpenStudy (anonymous):

a 3 foot pole costs a 6 foot shadow. a 4 foot pole casts an 8 foot shadow use a linear equation to predict the length of a shadow cast by a 6 foot pole. need help :)

OpenStudy (vera_ewing):

Ok, so shadow lengths are proportional to the heights of objects casting the shadows.

OpenStudy (vera_ewing):

Therefore, calling the shadow length l, the height h, and the proportionality constant k, l = kh. (The intercept is 0 because an object with no height casts no shadow.) Therefore, in this instance k = l/h = 6/3 or 8/4 = 2. then l(6) = 2 X 6 = 12 feet.

OpenStudy (anonymous):

ohhh.. I see I kinda got mixed up well thanks :)

OpenStudy (vera_ewing):

You're welcome!